# The Critical LYZ equation in Kähler Geometry

### v1

@@ 1.2. Our main results
Collins-Jacob-Yau [7] suggested studying the LYZ equation when $\theta\leq(n-2)\frac{\pi}{2}$. This was later explicitly posed as a question by Li [25], as resolving it is essential for the LYZ equation to induce a Bridgeland stability condition on $D^{b}\text{Coh}(M)$.
@@
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>\frac{n-3}{2}\pi.$ (5)
@@
Assume the same conditions as in Theorem 1.1 hold. Let $u^{t}$ be the smooth solution of (6) with $\sup_{M}u^{t}=0$. Then there exist uniform constants $C$ and $t_{0}$ depending only on $(M,\omega,\chi)$ but independent of $t$ such that for any $t\in(0,t_{0})$, we have
@@
According to Collins-Jacob-Yau [7], $\sup_{M}|u^{t}|$ is uniform bounded. Their results then imply that $C^{2,\alpha}$ estimates can be reduced to prove the uniform complex Hessian estimates. Our main contribution is the proof of the uniform complex Hessian estimates of Hou-Ma-Wu type
$\sup\limits_{M}|\sqrt{-1}\partial\bar{\partial}u|_{\omega}\leq C(1+\sup\limits_{M}|\partial u|^{2}_{\omega})$ (8)
and the uniform gradient estimates via the Dinew-Kolodziej blow-up argument [10].
However, the situation here is more complicated. If the gradient estimates fail, we obtain a $C^{1}$ bounded non-constant function $v$ on $\mathbb{C}^{n}$ which, in the sense of currents, is a weak solution to one of the following equations: $\sigma_{n}=0$, $\sigma_{n-1}=0$, or $\sigma_{n-1}+\sigma_{n}=0$. The last case is new and requires a new Liouville-type theorem. Note that $v+|z|^{2}$ is plurisubharmonic and $\Delta v$ is uniformly bounded in the weak sense (see Remark 4.2). When we follow the argument in [10, 36], we find the first case can be handled similarly. For the second case, we need to prove the standard mollification $v]_{r}$ of $v$ and $\frac{1}{3}[(v^{\epsilon})^{2}]_{r}-\tau_{0}|z|^{2}$ are subsolutions, where $v^{\epsilon}=[v]_{\epsilon}$ and $\tau_{0}>0$ is small enough. To address this issue, our starting point is that a smooth function $v$ satisfies equation $\sigma_{n-1}+\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n}$ is equivalent to the function $v+|z_{n+1}|^{2}$ satisfies equation $\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n+1}$, where $z_{n+1}$ is a new complex variable.
@@
Let $(M,\omega)$ be a compact Kähler manifolds with dimension $n=3$ and $\chi$ be a closed real $(1,1)$ form satisfying $\chi\wedge\omega>0$ as a positive $(2,2)$ form and
@@
Another consequence of our result is to solve the above Hessian quotient equation in dimension 4 under weaker conditions than the one by Sun [34] and Székelyhidi [36].
@@
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a closed real $(1,1)$-form satisfying $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a positive $(3,3)$ form}$ and
@@
The Hessian quotient equation $\sigma_{k}(\chi_{u})=c\sigma_{l}(\chi_{u})$ with $0\leq l<k\leq n$ was solved by Sun [34] and Székelyhidi [36] by different continuity methods.
The organization of this paper is as follows. In Section 2, we provide preliminaries on $k$-Hessian operators, the special Lagrangian operator $\theta_{\omega}$, and the approximating family $\theta(t)$. We also give the $C^{0}$-estimate for $u^{t}$ using results from [7]. In Section 3, we first proves a key inequality implied by the subsolution condition, then we establish the uniform complex Hessian estimate of Hou-Ma-Wu-type which is independent of $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$). In Section 4, based on the complex Hessian estimate, we prove the gradient estimate by a blow up argument. Then we prove the main theorems. In Section 5, we apply the main results to the 3-dimensional 2-Hessian equation and 4-dimensional Hessian quotient equation.
@@ 2.1. The k-Hessian operators and the special Lagrangian operator
For any vector $\lambda\in\mathbb{R}^{n}$ and $1\leq k\leq n$, we define
$\sigma_{k}(\lambda)=\sum\limits_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}.$ (10)
For a real $(1,1)$-form $\alpha$, the $k$-Hessian operator of $\alpha$ with respect to $\omega$ is defined by
@@
In local coordinates, let $\alpha=\sqrt{-1}\alpha_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}$ and $\omega=\sqrt{-1}g_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}$, we have $\sigma_{k}{(\alpha)}=\sigma_{k}(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}}))$. For simplicity, we also write $\sigma_{k}(g^{j\bar{l}}\alpha_{i\bar{l}})$ to represent $\sigma_{k}(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}}))$ .
@@
[…]
We have the concavity property for Hessian operators.
@@
The function ${\sigma_{n-1}^{\frac{1}{n-1}}}(\lambda)$ is concave in $\Gamma_{n-1}$. The function ${\sigma_{n-1}^{\frac{1}{n-1}}}(A)$ is concave in $\mathcal{M}(\Gamma_{n-1})$.
@@
$\theta(\lambda):=\sum_{i=1}^{n}\arctan\lambda_{i}\quad\textup{for}\ \lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$
and
$\Gamma^{\tau}:=\{\lambda\in\mathbb{R}^{n}\mid\theta(\lambda)>\tau\}\subset\mathbb{R}^{n}\quad\textup{for $\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$}.$
We have the following useful facts.
Lemma 2.2 (Yuan [42], Wang-Yuan [40]).
If $\theta(\lambda)\geq\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$ for $\lambda=(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, then we have
@@
    $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n-1}>0$, and $\lambda_{n-1}\geq|\lambda_{n}|$,
@@
Let $\hat{\theta}(t)$ be the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ and $\theta(t)=(n-2)\frac{\pi}{2}-\hat{\theta}(t)$. We first prove $\hat{\theta}(t)$ is sufficiently close to $\pi$ for sufficiently small $t>0$.
@@
If $\underline{u}$ is the subsolution of the critical LYZ equation (4), then there exists uniform positive constants $t_{0}$ and $C$ depending only on $\chi$ and $\omega$ such that $\hat{\theta}(t)\in(\pi-2Ct,\pi-Ct)$ for any $t\in(0,t_{0})$.
@@
By direct computation, we have
@@
Since
$\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}<0,$
for sufficiently small $t$, we have
$\mathrm{Re}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}<\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+Ct<0.$ (12)
Next we want to prove the imaginary part is positive. The subsolution condition implies the positivity
@@
Then for sufficiently small $t>0$, we get
$\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega=\mathrm{Im}\int_{M}(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}\wedge\omega>0$ (13)
Thus we get
@@
$+t\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
@@
where in the last inequality we use (13). Then there exists a uniform positive constant $C$ depending only on $\chi$ and $\omega$ such that the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ satisfies
$\hat{\theta}(t)=\mathrm{arccot}\Big(\frac{\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+O(t)}{t\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})}\Big)\in(\pi-2Ct,\pi-Ct).$
Hence by Lemma 2.3, we have, for sufficiently small $t>0$,
$\theta(t)\in\big((n-2)\frac{\pi}{2}+Ct,(n-2)\frac{\pi}{2}+2Ct\big),$ (14)
and then
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}}+t\omega)>(n-3)\frac{\pi}{2}+2Ct>\theta(t)-\frac{\pi}{2},$ (15)
which means that $\underline{u}$ is also a subsolution of the equations (6) for sufficiently small $t>0$. According to Lin [26] (see also Sun [35]), for every small $t>0$, there exists a smooth solution $u^{t}$ solving the supercritical LYZ equation (6). According to Proposition 3.6 in [7], we get the uniform $C^{0}$ estimates of $u^{t}$.
Lemma 2.4 (Collins-Jacob-Yau [7]).
Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$. Then there exist uniform constants $t_{0}$ and $C$ such that for any $t\in(0,t_{0})$, we have
$\sup_{M}|u^{t}|\leq C.$
@@ 3.1. The linearized operator
We first give some useful formulae for the complex Hessian estimate in [11].
@@
Let $u^{t}$ be the solution of the LYZ equation (6).
(i) The linearized operator $\mathcal{L}$ of the LYZ equation (6) has the form
$\mathcal{{L}}(v)=F^{i\bar{j}}v_{i\bar{j}},$
where $(F^{i\bar{j}})_{n\times n}$ is the inverse matrix of $wg^{-1}w+g$ with $g=(g_{i\bar{j}})_{n\times n}$, $w=(w_{i\bar{j}})_{n\times n}$ for $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u_{i\bar{j}}$.
(ii) Differentiating the LYZ equation (6), we have
$F^{i\bar{j}}w_{i\bar{j},p}=0,$ (16)
$F^{i\bar{j}}w_{i\bar{j},p\bar{p}}=F^{i\bar{l}}F^{k\bar{j}}w_{i\bar{j},p}(w_{k\bar{m},\bar{p}}g^{r\bar{m}}w_{r\bar{l}}+w_{k\bar{m}}g^{r\bar{m}}w_{r\bar{l},\bar{p}}),$ (17)
where we used covariant derivatives.
3.2. Key inequalities by the subsolution condition
The subsolution condition is very important to the $C^{2}$-estimates of solutions of fully nonlinear elliptic PDEs, one can see Guan [12] and Székelyhidi [36]. Similar to [12, 36], we prove key inequalities which will be used in the second order estimate. Fortunately, we can get a uniform constant $\delta_{0}$ in the following lemma.

### v2

@@ 1.2. Our main results
Collins-Jacob-Yau [7] suggested studying the LYZ equation when $\theta\leq(n-2)\frac{\pi}{2}$. This was later explicitly posed as a question by Li [26], as resolving it is essential for the LYZ equation to induce a Bridgeland stability condition on $D^{b}\text{Coh}(M)$, the bounded derived category of coherent sheaves on $M$.
@@
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>{(n-3)\frac{\pi}{2}}.$ (5)
@@
Assume the same conditions as in Theorem 1.1. Let $u^{t}$ be the smooth solution of (6) with $\sup_{M}u^{t}=0$. Then there exist uniform constants $C$ and $t_{0}$, depending only on $(M,\omega,\chi)$ but independent of $t$, such that for any $t\in(0,t_{0})$,
@@
According to Collins–Jacob–Yau [7], $\sup_{M}|u^{t}|$ is uniformly bounded. Their results imply that the $C^{2,\alpha}$ estimates can be reduced to proving uniform complex Hessian estimates. Our main contribution is the proof of the uniform complex Hessian estimate of Hou–Ma–Wu type,
$\sup\limits_{M}|\sqrt{-1}\partial\bar{\partial}u|_{\omega}\leq C\bigl(1+\sup\limits_{M}|\partial u|_{\omega}^{2}\bigr),$ (8)
and the uniform gradient estimate via the Dinew–Kolodziej blow-up argument [10].
However, the situation here is more subtle. If the gradient estimate fails, we obtain a bounded non-constant $C^{1}$ function $v$ on $\mathbb{C}^{n}$ that, in the sense of currents, is a weak solution to one of the following equations:
$\sigma_{n}=0,\quad\sigma_{n-1}=0,\quad\text{or}\quad\sigma_{n-1}+\sigma_{n}=0.$
The last case is new and requires a corresponding Liouville-type theorem. Note that $v+|z|^{2}$ is plurisubharmonic and $\Delta v$ is uniformly bounded in the weak sense (see Remark 4.2). When following the argument in [10, 38], we find that the first case can be handled similarly. For the second case, we need to show that the standard mollification $v]_{r}$ of $v$, as well as $\frac{1}{3}[(v^{\epsilon})^{2}]_{r}-\tau_{0}|z|^{2}$, are subsolutions, where $v^{\epsilon}=[v]_{\epsilon}$ and $\tau_{0}>0$ is sufficiently small. To address this issue, our starting observation is that a smooth function $v$ satisfies $\sigma_{n-1}+\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n}$ if and only if the function $v+|z_{n+1}|^{2}$ satisfies $\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n+1}$, where $z_{n+1}$ is a new complex variable. Based on this observation, we use potential theory to obtain the required Liouville-type theorem (see Theorem 4.1).
@@
Let $(M,\omega)$ be a compact Kähler manifolds with dimension $n=3$ and $\chi$ be a closed real $(1,1)$-form satisfying $\chi\wedge\omega>0$ as a positive $(2,2)$-form and
@@
Another consequence of our result is to solve the above Hessian quotient equation in dimension 4 under weaker conditions than those by Sun [36] and Székelyhidi [38].
@@
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a closed real $(1,1)$-form satisfying $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a positive $(3,3)$-form}$ and
@@
We remark that the Hessian quotient equation is of the form $\sigma_{k}(\chi_{u})=c\sigma_{l}(\chi_{u})$ with $0\leq l<k\leq n$. It reduces to the $J$-equation when $k=n$ and l=n1l=n-1 which was solved by Song-Weinkove [35] using the flow method, and the case $k=n,1\leq l\leq n-1$ was solved by Fang-Lai-Ma [11]. The general case was subsequently solved by Sun [36] and Székelyhidi [38] via different continuity methods.
The organization of this paper is as follows. In Section 2, we provide preliminaries on $k$-Hessian operators, the special Lagrangian operator $\theta_{\omega}$, and the approximating family $\theta(t)$. We also give the $C^{0}$-estimate for $u^{t}$ using results from [7]. In Section 3, we first prove a key inequality implied by the subsolution condition, then we establish the uniform complex Hessian estimate of Hou-Ma-Wu-type which is independent of $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$. In Section 4, based on the complex Hessian estimate, we prove the gradient estimate by a blow up argument. Then we prove the main theorems. In Section 5, we apply the main results to the 3-dimensional 2-Hessian equation and 4-dimensional Hessian quotient equation.
@@ 2.1. The $k$-Hessian operators and the special Lagrangian operator
For an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$ and $1\leq k\leq n$, we define
$\sigma_{k}(\lambda)=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}.$ (10)
For a real $(1,1)$-form $\alpha$, the $k$-Hessian operator of $\alpha$ with respect to the Kähler form $\omega$ is defined by
@@
In local coordinates, writing $\alpha=\sqrt{-1}\alpha_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}$ and $\omega=\sqrt{-1}g_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}$, we have
$\sigma_{k}(\alpha)=\sigma_{k}\!\bigl(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}})\bigr).$
For simplicity, we also denote $\sigma_{k}(g^{j\bar{l}}\alpha_{i\bar{l}})$ for $\sigma_{k}\!\bigl(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}})\bigr)$.
@@
$\mathcal{M}(\Gamma_{n-1})$ $:=\{A\mid\text{$A$ is an $n\times n$ Hermitian matrix and }\lambda(A)\in\Gamma_{n-1}\}.$
The Hessian operators possess the following concavity property.
@@
The function $\sigma_{n-1}^{\frac{1}{n-1}}(\lambda)$ is concave in $\Gamma_{n-1}$. Consequently, the function $\sigma^{\frac{1}{n-1}}_{n-1}(A)$ is concave in $\mathcal{M}(\Gamma_{n-1})$.
@@
$\theta(\lambda):=\sum_{i=1}^{n}\arctan\lambda_{i}\quad\text{for }\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n},$ (12)
and define the cones
$\Gamma^{\tau}:=\{\lambda\in\mathbb{R}^{n}\mid\theta(\lambda)>\tau\}\subset\mathbb{R}^{n}\quad\text{for }\tau\in\bigl[(n-2)\tfrac{\pi}{2},n\tfrac{\pi}{2}\bigr).$
The following lemma collects some key properties of these cones.
Lemma 2.2 (Yuan [44], Wang–Yuan [42]).
Assume that $\theta(\lambda)\geq\tau$ for some $\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$, and that the components of $\lambda=(\lambda_{1},\dots,\lambda_{n})$ satisfy $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$. Then the following hold:
@@
    $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n-1}>0$ and $\lambda_{n-1}\geq|\lambda_{n}|$;
@@
Let $\hat{\theta}(t)$ denote the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$, and set $\theta(t)=(n-2)\frac{\pi}{2}-\hat{\theta}(t)$. We first show that $\hat{\theta}(t)$ is close to $\pi$ when $t>0$ is sufficiently small.
@@
Assume that $\underline{u}$ is a subsolution of the critical LYZ equation (4). Then there exist uniform positive constants $t_{0}$ and $C$, depending only on $\chi$ and $\omega$, such that for all $t\in(0,t_{0})$,
$\hat{\theta}(t)\in\bigl(\pi-2Ct,\;\pi-Ct\bigr).$
@@
By a direct computation, we have
@@
Since $\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}<0$, it follows that for sufficiently small $t$,
$\mathrm{Re}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}<\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+Ct<0.$
We now show that the imaginary part is positive. The subsolution condition implies the positivity
@@
Hence, for sufficiently small $t>0$,
$\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega=\mathrm{Im}\int_{M}(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}\wedge\omega>0.$ (13)
Consequently,
@@
$\quad+t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
@@
where the last inequality uses (13).
Therefore, there exists a uniform constant $C>0$, depending only on $\chi$ and $\omega$, such that the principal argument $\hat{\theta}(t)$ of $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ satisfies
$\hat{\theta}(t)=\pi-\arctan\!\left(\frac{t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})}{-\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+O(t)}\right)\in(\pi-2Ct,\;\pi-Ct)$
for all sufficiently small $t>0$. ∎
Hence, for sufficiently small $t>0$, Lemma 2.3 yields
$\theta(t)\in\bigl((n-2)\tfrac{\pi}{2}+Ct,\;(n-2)\tfrac{\pi}{2}+2Ct\bigr),$ (14)
and consequently,
$\min_{1\leq j\leq n}\sum_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\arctan\lambda_{i}(\chi_{\underline{u}}+t\omega)>(n-3)\frac{\pi}{2}+2Ct>\theta(t)-\frac{\pi}{2}.$ (15)
This shows that $\underline{u}$ remains a subsolution of equation (6) for all sufficiently small $t>0$.
By the result of Lin [27] (see also Sun [37]), for each such $t$, there exists a smooth solution $u^{t}$ to the supercritical LYZ equation (6). Moreover, Proposition 3.6 of [7] provides a uniform $C^{0}$ estimate for these solutions.
Lemma 2.4 (Collins–Jacob–Yau [7]).
Let $\underline{u}$ be a subsolution of the critical LYZ equation (4), and let $u^{t}$ be the solution of (6) normalized by $\sup_{M}u^{t}=0$. Then there exist constants $t_{0}>0$ and $C>0$, depending only on the background data $(\chi,\omega)$ and $\underline{u}$, such that for all $t\in(0,t_{0})$,
$\|u^{t}\|_{C^{0}(M)}\leq C.$
@@ 3.1. The linearized operator
We first collect some useful formulas for the complex Hessian estimate from [12].
@@
Let $u^{t}$ be a solution of the LYZ equation (6). Then:
  1. (i)
    The linearized operator $\mathcal{L}$ of (6) takes the form
    $\mathcal{L}(v)=\sum_{i,j}F^{i\bar{j}}v_{i\bar{j}},$
    where the matrix $(F^{i\bar{j}})$ is the inverse of $wg^{-1}w+g$. Here $g=(g_{i\bar{j}})$ and $w=(w_{i\bar{j}})$ with $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$.
  2. (ii)
    Differentiating the LYZ equation (6) (using covariant derivatives) yields
    $\sum_{i,j}F^{i\bar{j}}w_{i\bar{j},p}=0,$ (16)
    $\sum_{i,j}F^{i\bar{j}}w_{i\bar{j},p\bar{p}}=\sum_{i,j,k,l}F^{i\bar{l}}F^{k\bar{j}}w_{i\bar{j},p}\bigl(w_{k\bar{m},\bar{p}}g^{r\bar{m}}w_{r\bar{l}}+w_{k\bar{m}}g^{r\bar{m}}w_{r\bar{l},\bar{p}}\bigr).$ (17)
3.2. Key inequalities from the subsolution condition
The subsolution condition plays a crucial role in the $C^{2}$-estimates for solutions of fully nonlinear elliptic PDEs, as demonstrated in the works of Guan [13] and Székelyhidi [38]. Following the approach in these references, we establish key inequalities that will be essential for the second-order estimate. A favorable feature is that the constant $\delta_{0}$ appearing below can be chosen uniformly.